OpenStudy (anonymous):

Hi, I need help with this problem: Sin(4x)+sin(6x) If you can show me step by step that would be much appreciated thank you.

OpenStudy (mathmate):

I assume you need to transform the sum into a product. Hints: Since 4x=5x-x, 6x=5x+x So Sin(4x)+sin(6x)=sin(5x-x)+sin(5x+x) Use sum and difference formula to simplify: sin(5x-x)=sin5x cosx-cos5x sinx sin(5x+x)=sin5x cosx+cos5x sinx Now you just have to add the two.